OpenStudy (anonymous):

Suppose R'(1)=16. a) If f(x)=R(5x), find f '(1/5) b) If g(x)=R(x+8), find g'(−7) c)If h(x) = R(x/7), find h'(7)

OpenStudy (anonymous):

@iambatman can you help me please?

jimthompson5910 (jim_thompson5910):

If f(x)=R(5x), then f'(x) = ??

OpenStudy (anonymous):

Right, I don't know if I am supposed to use the chain rule, but if so, what is the inside and what is the outside?

jimthompson5910 (jim_thompson5910):

the outer function is R(x) the inner function is 5x

OpenStudy (anonymous):

ok,so would it be R'(5x)*5? which would be 16*5=80

OpenStudy (anonymous):

because 5(1/5)=1 and R'(1)=16

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

oh wait, nvm

jimthompson5910 (jim_thompson5910):

f(x)=R(5x) f'(x)=5*R'(5x) f'(1/5)=5*R'(5(1/5)) f'(1/5)=5*R'(1) f'(1/5)=5*16 f'(1/5)=80 yeah you got it right

OpenStudy (anonymous):

ok, and so for b), would the answer be -112 because after it is all simplified, you get 16*-7

jimthompson5910 (jim_thompson5910):

g(x)=R(x+8) g'(x)=R'(x+8) g'(-7)=R'(-7+8) g'(-7)=R'(1) g'(-7)=16 so this shows you that shifting the graph (or translating it) does not affect the slopes of the tangent lines.

jimthompson5910 (jim_thompson5910):

I didn't show it, but with the chain rule you'll have 1* on the outside, but that just results in R'(x+8)

OpenStudy (anonymous):

ok, I see what I did wrong. I put it as R'(-7+8)*(x) which gave me R'(1)=16*-7. however the derivative of x is 1 which I messed up.

OpenStudy (anonymous):

I go the answer for c) by doing the chain and quotient rule. the answer I got was 16/7

jimthompson5910 (jim_thompson5910):

no you would derive x+8 to get 1 and then multiply it by R'(x+8)

jimthompson5910 (jim_thompson5910):

h(x) = R(x/7) h'(x) = (1/7)*R'(x/7) h'(7) = (1/7)*R'(7/7) h'(7) = (1/7)*R'(1) h'(7) = (1/7)*16 h'(7) = 16/7 looks good

jimthompson5910 (jim_thompson5910):

you don't need to do the quotient rule, but it's possible to use it

jimthompson5910 (jim_thompson5910):

I'd just pull out the coefficient

OpenStudy (anonymous):

ok, thanks!! I just have one last question to do for my homework. Can you also help me with that really quick?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

I need to find the derivative of g(x)=5/(a^2-w^2)^4 My answer was -40a(a^2-w^2)^3/(a^2-w^2)^8

OpenStudy (anonymous):

\[\frac{ -40a(a^2-w^2)^3 }{ (a^2-w^2)^8 }\] to make my answer more clear.

jimthompson5910 (jim_thompson5910):

I see g(x), but I don't see any x on the right side

OpenStudy (anonymous):

sorry I meant to put g(w) but habit forced me to put x

jimthompson5910 (jim_thompson5910):

how did you get -40a? When you derive the inside a^2 - w^2, you should be left with 2w (the a^2 is constant which derives to 0)

jimthompson5910 (jim_thompson5910):

-2w I mean

OpenStudy (anonymous):

now that you mention that, I have no clue because I got -40a by -5*8a and I got 8a by 4*2a

jimthompson5910 (jim_thompson5910):

it's very close to the answer, but incorrect

jimthompson5910 (jim_thompson5910):

you can't say a^2 turns into 2a because you aren't deriving with respect to 'a'

OpenStudy (anonymous):

right, I always get confused when there are just letters...so would the new answer be -10w(a^2-w^2)^3/(a^2-w^2)^8

jimthompson5910 (jim_thompson5910):

the general rule is to derive with respect to the independent variable

jimthompson5910 (jim_thompson5910):

usually (not always), the other letters are constants

OpenStudy (anonymous):

right, I get that, it just always trips me up...

jimthompson5910 (jim_thompson5910):

-10w(a^2-w^2)^3/(a^2-w^2)^8 is close, but again, incorrect

OpenStudy (anonymous):

what part of it is incorrect?

jimthompson5910 (jim_thompson5910):

it might help to think of the problem in the form u/v

OpenStudy (anonymous):

the -10w, the (a^2-w^2)^3, or the bottom?

OpenStudy (anonymous):

what do you mean u/v?

jimthompson5910 (jim_thompson5910):

u = 5 v = (a^2-w^2)^4

jimthompson5910 (jim_thompson5910):

so u' = 0 v' = 4(a^2-w^2)*2w = 8w(a^2-w^2)

jimthompson5910 (jim_thompson5910):

this is setting up for the quotient rule

jimthompson5910 (jim_thompson5910):

oops, v' = 4(a^2-w^2)*(-2w) = -8w(a^2-w^2), there fixed it

OpenStudy (anonymous):

ok, so it would be -40w.....

OpenStudy (anonymous):

first off, is the bottom part of my answer correct?

jimthompson5910 (jim_thompson5910):

oh and I completely forgot an exponent v' = 4(a^2-w^2)^3*(-2w) = -8w(a^2-w^2)^3

OpenStudy (anonymous):

or wait is it just 40w since it is -5*-8w...

jimthompson5910 (jim_thompson5910):

anyways, u'*v - v'*u = 0*(a^2-w^2)^4 - ( -8w(a^2-w^2)^3 )*5 = 40w(a^2-w^2)^3 is all in the numerator

OpenStudy (anonymous):

ok, thanks!! I think it have it now. \[\frac{ 40w(a^2-w^2)^3 }{ (a^2-w^2)^8 }\]

jimthompson5910 (jim_thompson5910):

then you can reduce to get \[\large \frac{ 40w(a^2-w^2)^3 }{ (a^2-w^2)^8 } = \frac{40w}{(a^2-w^2)^5 } \]

OpenStudy (anonymous):

ok, thanks!!! I finally get it!!

jimthompson5910 (jim_thompson5910):

np